OpenStudy (anonymous):

Mary and her brother John collect foreign coins. Mary has four times the number of coins that John has. Together they have 150 foreign coins. Find how many coins Mary has.

OpenStudy (anonymous):

Let the number of coins John has be "x". The the number of coins Mary has is "4x" ------------- EQUATION: x + 4x = 150 5x = 150 x = 30 John has x = 30 coins Mary has 4x = 4*30 = 120 coins.

OpenStudy (anonymous):

J + 4J = 150 J = 150/5 = 30 coins -> M has 4* 30 = 120 coins